OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1)with a slope of 4 y - 1 = 4(x - 3) y - 3 = 4(x + 1) y + 1 = 4(x + 3) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y+1=4(x+3)

hero (hero):

\[\text{The point slope form is} \space y-y_1 = m(x-x_1)\]